OpenStudy (anonymous):

A piece of lead weighs 0.2kg. It takes 300 seconds for it to reach melting point of 600 degrees C from an initial temperature of 0 degrees C. The specific heat capacity of lead is 130 J kg^-1 k^-1. Calculate the total energy supplied to the lead before it starts to melt.

OpenStudy (anonymous):

melting point of lead is 600 K and not °C

OpenStudy (anonymous):

\[\Delta H=\int\limits_{T1}^{T2} cp dT\]

OpenStudy (anonymous):

DH= 130 Jkg-1K-1 * (600-0) = 78000 Jkg-1 DH= 78 kJkg-1 * 0,2kg = 15,6 kJ

OpenStudy (anonymous):

and q= DH/t= 15,6 kJ/300s = 52 W

OpenStudy (anonymous):

Is DH the same as U = m c delta T?

OpenStudy (anonymous):

well yes, i dont know what letter you use to mark it...

OpenStudy (anonymous):

Ok. For the first part of the question. Q is the energy supplied or removed. M = mass in kg, C = specific heat capacity and Delta T is the change in temperature.

OpenStudy (anonymous):

Could you explain it that way. I do not understand your terminology.

OpenStudy (anonymous):

what do you need explained?

OpenStudy (anonymous):

DH= 130 Jkg-1K-1 * (600-0) = 78000 Jkg-1. Is this the answer to the first part of the question. I have done this calculation. U = m c delta T. 0.2*130Jkg-1 K-1 *600K. Is that correct?

OpenStudy (anonymous):

yes it is correct and if you compare it i did the same

OpenStudy (anonymous):

Ok great. Now what about the second part of question when i ask for the rate of heating of the lead?

OpenStudy (anonymous):

well taht is q in W and W is J/s

OpenStudy (anonymous):

heat flow rate

OpenStudy (anonymous):

I have'nt done any work on heat flow rate to be honest!

OpenStudy (anonymous):

well then it's time to start! :)

OpenStudy (anonymous):

(1) Q = kA delta T / d ?

OpenStudy (anonymous):

well that is different equation...

OpenStudy (anonymous):

What is the formula for heat flow rate ?

OpenStudy (anonymous):

dont you have your textbooks or scripts for studying?

OpenStudy (anonymous):

Yes i have advanced physics for you. But i cannot find any formula for the heat flow rate in it.

OpenStudy (anonymous):

ok lets think about that for a second, how do you mark heat? lets say Q and how do you mark time, lets say t and rate of heat transfer is Q/t and lets say we mark that with h, or q and there you have it...

OpenStudy (anonymous):

i must admit that its been about 5 years i studied that, nowdays i just get DSC or DMA or FTIR or some other method diagrams and dont have to deal with that...

OpenStudy (anonymous):

Ok. Q is energy supplied or removed / T = change in temperature (rise and fall) ?

OpenStudy (anonymous):

you take heat and take its change in time and you get amount of heat given to the body or heat that body releases in time...

OpenStudy (anonymous):

it is like speed generaly, so if "s" is distance and "t" is time so s/t is speed of travel...

OpenStudy (anonymous):

So it will be Q / T = ?. Is that it

OpenStudy (anonymous):

yes but time is usually marked with small t

OpenStudy (anonymous):

Ok. So it will be Q/ delta t = 15600J / 300s = 52J ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and it is J/s or J s-1 or W

OpenStudy (anonymous):

Ok thanks it looks more simple than it should be!